# Another look at inference after prediction

### v3

@@ 5.2 Results
Figure 2 displays the CI length ratio across the different settings. In Scenario 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ perform similarly, with CI length ratios below 1, indicating that the PB inference methods have reduced standard error relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Comparing Scenario 1a and 1b, we note a significant decrease in the gains in statistical efficiency of the PB inference methods due to the weaker prediction model used in Scenario 1b. However, all methods are at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, as the sufficient condition in Corollary 1.1 is satisfied. Additionally, as the labeled set size increases and hence $\pi=n_{\text{\tiny{lab}}}/n$ increases, the efficiency gain decreases as expected. $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is always more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, particularly when the labeled dataset size is larger, due to the difference in the augmentation terms of these two estimators. In this scenario, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$.
Scenario 2 highlights the differences in performance between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, which underscore our findings in Corollary 1.1 and Corollary 2.1. In both sub-scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ consistently outperforms both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In Scenario 2a, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves an average CI length ratio of 0.9 across the various sample sizes, indicating a gain in efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In Scenario 2b, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is comparable to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for the larger labeled dataset sizes as there is not much to gain from the unlabeled data due to its relatively small size and the fact that $\widehat{Y}$ is a weak prediction of $Y$. In contrast, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ are outperformed by $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ in nearly all scenarios. In Scenario 2a, these estimators have CIs up to 30% longer than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for the larger labeled dataset sizes. This behavior is exacerbated in Scenario 2b, where the CI of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is up to 50% longer than the CI for $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and up to 72% longer for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$.
Labeled set size
Method 1000 2000 3000 4000 5000
Scenario 1a
     $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$0.181 0.124 0.650 0.430 0.165
     $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ 0.035 -0.042 0.015 0.005 -0.005
     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.004 0.010 -0.004 -0.018 -0.001
     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ 0.007 0.000 0.002 -0.009 -0.003
     $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ 0.007 0.000 0.002 -0.009 -0.003
Scenario 1b
     $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$1.070 0.181 0.426 -0.423 -1.205
     $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ 0.008 0.002 -0.016 -0.001 0.004
     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.007 -0.001 -0.003 -0.008 -0.010
     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ 0.007 $0$0.007 -0.006 -0.003
     $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ 0.006 $0$0.007 -0.006 -0.003
Scenario 2a
     $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$46.980 -47.523 -45.946 -46.725 -47.695
     $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$0.033 -0.023 0.027 0.004 0.005
     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$0.023 -0.009 -0.003 -0.032 -0.003
     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$0.024 -0.012 0.006 -0.018 0.001
     $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$0.030 -0.016 0.016 -0.007 0.003
Scenario 2b
     $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$64.272 -64.115 -63.307 -63.759 -64.627
     $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$0.012 -0.003 0.002 0.012 0.013
     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.000 0.001 -0.006 0.012 0.009
     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$0.001 $0$0.004 0.012 0.010
     $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$0.014 -0.006 -0.003 0.011 0.010
Table 1: Percent biases of the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$), the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$), the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$), and the naive estimator ($\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$). Negative values indicate underestimation, while positive values indicate overestimation of the true parameter.
@@
Figure 2: Confidence interval length ratio of the PB inference methods compared to baseline estimation with the labeled data ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$). A ratio less than 1 indicates that the PB inference method has a confidence interval that is shorter than the confidence interval using $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
@@ 6 UK Biobank Analysis
Central obesity, defined as excessive accumulation of body fat around abdominal region, is strongly associated with cardiometabolic diseases such as type 2 diabetes and cardiovascular disease. A major challenge in studying central obesity is the limited availability of direct body composition measurements. Android fat mass (AFM), measured via a dual-energy X-ray absorptiometry (DEXA) scan, is a key marker for assessing central obesity in clinical practice [Zhu et al., 2022]. However, due to its high cost, DEXA-derived AFM is usually only measured in a small subset of participants in large-scale studies. For example, in the UK Biobank, a prospective cohort study measuring genetic, lifestyle and health information on approximately 500,000 UK participants, DEXA-derived AFM is incompletely ascertained, particularly among populations populations historically underrepresented in genetic research [Sudlow et al., 2015]. For instance, out of $5,397$ South Asian participants in the UK Biobank, only $288$ (5.3%) participants have undergone DEXA scans.
[…]
@@
Table 2 presents point estimates, standard errors, and CI ratios (or equivalently standard errors length ratios) across the two prediction model settings for four estimators: the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$), the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$), and the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$). Echoing our theoretical and simulation study results, all methods yield similar point estimates as they are all consistent estimators for the same parameter. For the analysis with the strong prediction model, $\widehat{Y}$ provides additional predictive power beyond the regression of $Y$ on ${\bf{X}}$, satisfying the sufficient condition in Corollary 2.1. Consequently, both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In fact, all three PB inference methods ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$) perform similarly, yielding substantially smaller standard errors compared to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with average CI ratios below 0.85 for all methods. The analysis with the weak prediction model highlights the difference in performance between the PPI and CC methods and reinforces the findings in Corollary 2.1 and Corollary 2.2. Here, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ yield larger standard errors than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for all parameters of interest, with CIs 38.5% longer on average. In contrast, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is no worse than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for all parameters of interest.
Strong Prediction Model
Intercept Age Sex BMI
$\hat{\bm{\beta}}^{\text{\tiny{lab}}}$
EST -4.857 -0.003 0.813 0.173
SE 0.308 0.004 0.070 0.009
CI Ratio 1.00 1.00 1.00 1.00
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$
EST -4.513 -0.007 0.904 0.164
SE 0.261 0.004 0.063 0.007
CI Ratio 0.85 0.86 0.89 0.80
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$
EST -4.527 -0.007 0.900 0.165
SE 0.260 0.004 0.063 0.007
CI Ratio 0.85 0.86 0.89 0.80
$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$
EST -4.525 -0.006 0.905 0.163
SE 0.255 0.004 0.062 0.007
CI Ratio 0.83 0.85 0.89 0.77
Weak Prediction Model
Intercept Age Sex BMI
$\hat{\bm{\beta}}^{\text{\tiny{lab}}}$
EST -4.857 -0.003 0.813 0.173
SE 0.308 0.004 0.070 0.009
CI Ratio 1.00 1.00 1.00 1.00
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$
EST -4.530 -0.006 0.835 0.165
SE 0.416 0.006 0.098 0.012
CI Ratio 1.35 1.43 1.40 1.36
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$
EST -4.546 -0.006 0.834 0.165
SE 0.415 0.006 0.098 0.012
CI Ratio 1.35 1.43 1.40 1.36
$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$
EST -4.779 -0.004 0.819 0.171
SE 0.294 0.004 0.072 0.008
CI Ratio 0.95 0.96 1.03 0.93
Table 2: UK Biobank analysis of android fat mass based on the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$), the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$), the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$). The strong prediction model has an $R^{2}$ of 0.8, while the weak prediction model has an $R^{2}$ of 0.5. Estimates (EST), standard errors (SE), and confidence interval (CI) ratios are reported to evaluate performance. A CI ratio less than 1 indicates that a method yields a confidence interval shorter than the CI using $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
@@ 7 Discussion
In this paper, we showed that adding a simple weight to the PPI estimator provides valid PB inference that is guaranteed to be at least as efficient as estimation using only labeled data. The weight is drawn from the work of Chen and Chen [2000] in the context of double sampling. In contrast, the original PPI estimator can be less efficient than estimation using only the labeled data, a property that threatens its practical utility. By incorporating the proposed weight, the augmented estimation approach principally balances the contribution of labeled and unlabeled data and extends the applicability of the PPI method to scenarios where leveraging all available data is essential for making valid and efficient inference. In addition to theoretical results that support these findings, our simulation studies and application to UK Biobank data demonstrate that the PPI method can result in substantial efficiency loss when the ML model is inaccurate. Additionally, we provide connections to several existing approaches from the economics and statistics literature that provide interesting context for the PB inference problem. While we focused on least squares estimation for exposition, our results extend to other settings. The work of Chen and Chen [2000] and Angelopoulos et al. [2023a] apply to a wide variety of settings beyond linear regression, including Z-estimation problems.
The work of Angelopoulos et al. [2023a] was pivotal in advancing the growing body of literature in PB inference. Recent work has considered settings where the ML model is not black box and where active learning is utilized to obtain the labeled data [Zrnic and Candès, 2024b, b]. Additionally, two concurrent pre-prints have noted potential losses in efficiency of PB inference approaches. The authors of the PPI paper introduced a new method, PPI++, which aims to address this limitation of the PPI estimator [Angelopoulos et al., 2023b]. The PPI++ method is designed to adaptively minimize the trace of the asymptotic variance-covariance of the proposed estimator to ensure that it yields an improvement over an approach based on labeled data. In contrast, the approach of Chen and Chen [2000] minimizes the mean square error among a class of estimators with a linear augmentation term. Simultaneously, Miao et al. [2023] introduced the PoSt-Prediction Adaptive inference (PSPA) estimator that is both robust to the accuracy of the ML model and guaranteed to be at least as efficient as estimation using only labeled data. Unlike the PPI++ estimator, the PSPA estimator is designed to achieve element-wise variance reduction and was shown to outperform the PPI++ approach. The PSPA approach is based on augmenting estimating equations with information from $\widehat{Y}$, while the CC method provides a simple and natural extension to the PPI method by adding a weight to the linear augmentation term.
@@
Title:
The supplementary materials include proofs for theorems and additional examples for Corollary 1.1 (appendix.pdf).

### v4

@@ 5.2 Results
Figure 2 displays the relative efficiency across the different settings. In Scenario 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ perform similarly, with RE significantly above 1, indicating that the PB inference methods have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Comparing Scenario 1a and 1b, we note a significant decrease in the gains in statistical efficiency of the PB inference methods due to the weaker prediction model used in Scenario 1b. However, all methods are at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, as the sufficient condition in Corollary 1.1 is satisfied. Additionally, as the labeled set size increases and hence $n_{\text{\tiny{lab}}}/n$ increases, the efficiency gain decreases as expected. $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is always at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ in Scenarios 1a and 1b, and more efficient in Scenario 1a when the labeled dataset size is larger due to the difference in the augmentation terms of these two estimators.
Scenario 2 highlights the performance differences between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, which underscore our findings in Corollary 1.1 and Corollary 2.1. In both sub-scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ consistently outperforms both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ and is substantially more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In Scenario 2a, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves an average RE of 1.27 across the various sample sizes, indicating a significant gain over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In Scenario 2b, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ again demonstrates superior RE compared to all other methods and remains more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ offers a modest efficiency gain in Scenario 2a and comparable performance to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ in Scenario 2b. Both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ are outperformed by $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ in nearly all scenarios. In Scenario 2a, while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ remains comparable to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, the RE for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is as low as 0.88 with $n_{\text{\tiny{lab}}}=2000$. Both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ demonstrate a substantial loses in efficiency in Scenario 2b, with $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ exhibiting a RE of 0.63 with $n_{\text{\tiny{lab}}}=2000$.
@@
Figure 2: Relative efficiency of the PB inference methods compared to baseline estimation with the labeled data ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$). A ratio greater than 1 indicates that the PB inference method has a smaller mean squared error (MSE) than that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
@@ 6 UK Biobank Analysis
We also evaluated the performance of the PB inference methods in an analysis focused on central obesity. Central obesity is excessive accumulation of body fat around the abdominal region and is strongly associated with cardiometabolic diseases such as type 2 diabetes and cardiovascular disease. A major challenge in its study is the limited availability of direct body composition measurements. Android fat mass (AFM), measured via a dual-energy X-ray absorptiometry (DEXA) scan, is a key marker for assessing central obesity in clinical practice. 51 However, due to its high cost, DEXA-derived AFM is usually only measured in a small subset of participants in large-scale studies. For example, in the UK Biobank,52 a prospective cohort study measuring genetic, lifestyle, and health information on approximately 500,000 participants in the UK, DEXA-derived AFM is incompletely ascertained, particularly among populations historically underrepresented in genetic research. 52 For instance, out of $5,397$ South Asian participants in the UK Biobank, only $288$ (5.3%) participants have undergone DEXA scans.
Our goal is to examine associations between AFM and age, sex, and body mass index (BMI) among the South Asian population. Previous research has indicated that this population is at higher risk of central obesity and that it is associated with cardiometabolic diseases at lower BMI thresholds than European populations. 53 Since DEXA-derived AFM is missing for most South Asian participants, it provides a natural opportunity to leverage ML-predicted AFM in downstream analyses. To illustrate the impact of different ML models on the efficiency of the PB inference methods, we trained two random forest models in the White British population ($32,460$ participants) and used them to predict missing AFM values in the South Asian population. We built a “strong” prediction model that incorporated both the covariates of interest (i.e., age, sex, and BMI) as well as additional predictors, including smoking status, alcohol consumption, two bioelectrical impedance measurements of the legs, physical activity scores from the International Physical Activity Questionnaire, and sedentary behavior quantified as total hours spent driving, using a computer, and watching television. We built another “weak” prediction model that included only the additional predictors. The strong prediction model achieved an $R^{2}$ of 0.8 while the weak prediction model had a substantially lower $R^{2}$ of 0.5 in the South Asian subpopulation.
@@
Table 3 presents point estimates, standard errors, and RE, defined as the ratio of the variance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ to the variance of the given estimator, across the two prediction model settings for five estimators: the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$), the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$), the PSPA estimator ($\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$), and the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$). We do not include naive inference in our analyses due to the limitations demonstrated in our simulation studies. Echoing our theoretical and simulation study results, all methods yield similar point estimates as they are all consistent estimators of the same parameter. In the analysis with the strong prediction model, the predictions provide information on central obesity beyond what is captured by the covariates of interest, satisfying the sufficient condition in Corollary 2.1. Consequently, both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In fact, all four PB inference methods ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$) yield substantially smaller standard errors compared to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Among the four PB inference methods, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ show comparable performance, with average REs of 1.39, 1.39, and 1.47, respectively. $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ exhibits higher efficiency gains with an average RE of 1.68. The analysis with the weak prediction model highlights the differences in performance between the PPI and CC methods and reinforces the findings in Corollary 2.1 and Corollary 2.2. Here, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ yield larger standard errors than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for all parameters of interest, with an average RE of 0.52. In contrast, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, as well as $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$, outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, with average RE values of 1.28 and 1.16, respectively. Overall, the CC estimator is the best performing method in terms of efficiency, underscoring its utility in practice.
@@ 7 Discussion
In this paper, we bridge classical statistical methods with a modern problem to equip practitioners and methodologists with a rigorous approach for PB inference. We find that incorporating a simple weight in the PPI estimator previously proposed for studies with double-sampling designs22 yields valid inference that is guaranteed to be at least as efficient as estimation using only labeled data. Our work addresses efficiency shortcomings of the original PPI estimator which can undermine its practical value. Alongside theoretical results supporting these findings, our simulation studies and application to UK Biobank data verify that the PPI method can suffer substantial efficiency loss when the ML model is inaccurate. Additionally, we draw important connections to several established approaches from the economics and statistics literature to provide context for ongoing developments in PB inference.
Our work is not without limitations. First, we focused our analyses on least squares estimation following Motwani and Witten.20 Further work is needed to analyze other statistical procedures (e.g., Z-estimation problems in general). Second, we assumed that the labeled and unlabeled data are drawn from the same underlying distribution or equivalently, that $Y$ is missing completely at random. This assumption may not hold in practice, particularly when samples cannot be randomly selected for labeling due to study design constraints or the use of alternative sampling strategies to enhance efficiency (e.g., stratified sampling). However, the CC estimator can be extended to settings where $Y$ is missing at random by incorporating inverse probability weighting and regression adjustment. Lastly, we focused on the traditional missing data scenario where the proportion of labeled data, $n_{\text{\tiny{lab}}}/n$, is bounded away from 0. Analyses of semi-supervised settings where $\lim_{n_{\text{\tiny{lab}}}\to\infty}n_{\text{\tiny{lab}}}/n=0$ would be of interest to formally model scenarios where the size of labeled and unlabeled are of different orders of magnitude.
@@
The supplementary materials include proofs for theorems and additional examples for Corollary 1.1.